Your spaceship lands on an unknown planet. To determine the characteristics of this planet, you drop a wrench from 7.00 mm above the ground and measure that it hits the ground 0.812 ss later.

Part A.

What is the acceleration of gravity near the surface of this planet?

Express your answer in meters per second squared to three significant figures.

Part B.

Assuming that the planet has the same density as that of earth (5500 kg/m3kg/m3), what is the radius of the planet?
Answer #1

__________________________

(a)

Using h = ut + at^2 /2

7 = 0 + a (0.812^2) /2

a = 21.233 m/s^2

__________________________

(b)

a = GM / R^2

M = mass = volume x density = (4 x pi x R^3 / 3) x 5500

a = (6.67 x 10^-11 x 4 x pi x R x 5500) / 3 = 21.233

R = 1.3817798158 x 107 m

__________________________

Please rate

if any mistake in this answer please comment i will clarify your doubt . thank you